The quantity represented by is a function that changes over time (i.e., is not constant).
Part A
The quantity represented by is a function that changes over time (i.e., is not constant).
true
false





Part B
The quantity represented by is a function of time (i.e., is not constant).
true
false





Part C
The quantity represented by is a function of time (i.e., is not constant).
true
false





Part D
The quantity represented by is a function of time (i.e., is not constant).
true
false





Part E
A particle moves with constant acceleration . The expression represents the particle's velocity at what instant in time?
at time
at the "initial" time
when a time has passed since the beginning of the particle's motion, when its velocity was





More generally, the equations of motion can be written as


and
.

Here is the time that has elapsed since the beginning of the particle's motion, that is, , where is the current time and is the time at which we start measuring the particle's motion. The terms and are, respectively, the position and velocity at . As you can now see, the equations given at the beginning of this problem correspond to the case , which is a convenient choice if there is only one particle of interest.
Answer #1



Equations missing in the given question. Based on standard question in the book, I am solving this question as follows:


Kinmatic eqautions:


x(t)=xo + Vot + (1/2)at^2
   v(t)= Vo + at

   x(t) is the position of the particle at a given instant of time
    xo is the initial position
    v(t) is the velocity of the particle at a given instant of time
    vo is the intial velocity of the particle
     a is the acceleration


(A) The quantity represented by x is a function that changes over time (i.e., is not constant).


TRUE (as the equation includes the term t)

(B) The quantity represented by xo is a function of time (i.e., is not constant).


FALSE (as xo is independent of t )

(C) The quantity represented by Vo is a function of time (i.e., is not constant).


FALSE (as initial velocity is independent of t)

   (D) The quantity represented by v is a function of time (i.e., is not constant).


TRUE (as the equation includes t)


   (E) A particle moves with constant acceleration . The expression ( v(t)= Vo + at) represents the particle's velocity at what instant in time.
    
     Ans: when a time has passed since the beginning of the particle's motion.
Answer #2
Okay these are very easy and basic physics questions. Part A True Part B False Part C False Part D True Part E V^2=V^2_i+2n(x-x_i) The third answer Part F when a time t has passed since the particle's velocity was V_i The third answer. Part G x_B(t)=x_i + 0.5v_i(t-t_1)+a(t-t1)^2 The sixth answer Part H t=2t_1+V_i/2a The second answer I hope that this helps! Engineer.Extraordinaire